OpenStudy (anonymous):

suppose that y varies inversely with x. write a function that models the inverse function. x =1 when y = 12

OpenStudy (anonymous):

would it be y = 12/x or y=1/12x

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

???

OpenStudy (jdoe0001):

keep in mind that the "constant of variation" for an INVERSE variation will x* y so y = n/x n = x* y when x = 1, y = 12, so n = 1* 12 => 12 so the function will be y = 12/x